Tiedown mods/failures

28jul03

Log of major structural changes:
This a log of tiedown failures and modifications.
 
 
Log of all mods/failures
date 
yymmdd

td description
171115

8
replaced cable ldcell td8-2 (broken during maria). td8-1 still won't balance the -200mv offset. set values to always read 0.
170901

8
td 8 jack replaced
170829

4
td 4 ldcell 4-2 fails
170827

12
td 12 replaced
170825

4
12
td 4 replaced.. it had locked up..(more info)
started to hear noises in 12.. looks like it has a problem with a bearing.
1606xx

12
load cell 12-1 fails after 1606 hf run.
160317

8
new load cell connected 8-2.. need full calibration after painting done (and we can move the az, dome).
160113

8
finally tried to recalibrate load cell td8-2 wouldn't balance. looks like the arcing back on 151110 damaged the load cell. It will be replaced.
151110

8
cable for load cell td8-2 had arcing through outer protective sleeve (close to yoke). this was during the hf run.
cable eventually replaced
150914

4
watch dog time was releasing for a few seconds every now and then, and then running ok
don't know whether it was watchdog timer or little start. replaced watchdog timer
For debugging:
  • If you see the velocity feedback always going down to about 300.. this is probably the watchdog timer causing the value.
  • the motor speed monitor runs at about 400 counts
  • The watchdog always uses the same drive  to amplifier...
140923

12,8
lightning strike damaged td 12,td8
td12 came back on line later that afternoon (replaced little star)
td8 remained offline till .. 28sep14.. trouble with 485 communications.
we manually moved td8 to 12.5 inches on 26sep14 so platform closer to focus at night.
130302

4
tiedown 4  position jumps to 24 inches, stays there awhile, then jumps back.
120702

4
changed motor on td 4.
120516

8
felipe soberal writes:
During the inspection and lubrication process performed at Platform Tie Down #8 last Monday, it was identified that we also have a worn ball screw on the actuator jack.  Approximately three tooth on the lower end of the screw when the jack is fully extended were found broken.  This jack has been in service since June 2, 2005.
120423

12
jack replaced.
120419

12
td12 stops with  foldback, cmdspderr errors. Td12 jack making loud noises when pulling down. Torque measured in td12 much larger than the other tiedowns, for similar kips.
Ended up having to replace the jack.
120418

12
one of the new little stars (from the large order) installed in td 12 for testing.
120227

8
recalibrated td 8 load cell.
120225

12,4,8
    • Prior to 25feb12 the 1 sec tick was sent down on a single fiber to box 12-7. It was then split  and sent to the individual tiedowns. The splitter was lossy (about 13 db). The  1 sec tick rcvrs were receiving power levels near -25dbm. This was lower than the required power level (-24dbm)  for the FO rcvr chip.
  • On 25feb12 a fiber optic buffer was installed in the back of the pnt vme crate. It takes a single fiber input and generates 3 separate fiber outputs. These are then sent to tiedowns 12,4,8 on separate fibers (so the splitter is no longer needed). The table  shows the optical power levels measured on 26feb12 after the new 1 sec tick fibers were installed.
120222

12,4,8
  • Fiber optic cables td12  from fiber box to huffman box replaced.
  • optical power splitter for 1 sec tick at td12. tick enters from upstairs, splits into 4 outputs.
  • Used optical xmt,power meter to measure  loss of this splitter. The are the values from the tester xmiter thru device to power meter measurement. It shows the loss in the splitter (but these are not the actual values used by the 1 sec tick xmt/rcvr).
    • 850 um
    • input power: -14dbm
    • output power:
      • to td12 : -27.4dbm
      • to td8   : -24.7 dbm
      • to td4   : -27.3 dbm
      • spare    : -27.3 dbm
  • switched 1 sec tick spare to got to td4 (to see if anything got better).
120219

4,8
change of 120217 did not fix the drifting of td4 clock.
at 10:42 on 120219 i switched the order that the vxWorks program talks to td4,td8
before : td12,td4,td8
now    : td12,td8,td4
see if the time drifts switches to dt8...(120220... drift remained at td4).

120217

4
updated software to declare tmscounter tmscounteratlast tick to be atomic  accesses (updated in the afternoon)
120209

12,4,8
Have been getting cmdspdErr after switching to the new operating system. (cmdVel - motorSpd > 50 cnts).
-Measured a/d for spdMon in lab. 46.8K resistor way to big to map 8 volts to 2.5 volts.
- switched  46.8K resistor to 22k got closer to correct value but still got error.
-Looking back in old data, this difference was much greater than with the 22K resistor. but the general status
  bit saying cmdSpdErr was never set in the data read back from the little star.
- There must have been a bug in the previous version of the software so that the status bit was not getting set. The new compiler fixed it?
- The rp4 resistors being removed on the adc boards got rid of  the abilitiy to have a reliable voltage divider.
120208

4,8
switched little stars to see if td4 running fast stays at 4 or moves to 8.
(problem remained at 4)
120124

12,4,8
2 new little stars installed in td4,8, all 3 little stars now running program compiled with version 6. of dynamic C.
111108

4,8
td 4,8 little stars offline.. for rest of the year.
111031

4
changed dac signal conditioning board to see if this helps the problem from 27,29oct11.
111029

4
tiedown 4 made load noises again at 5:21 am.
111027

4
tiedown 4 made loud noise at 1:17 am
11713

12,8,4
went offline 14:16 13jul11.
td12 encoder reading bad
4,8 no communication.
4,8 remain offline for a few weeks.
110609

12,8
recalibrated load cells 12,8. td12-2 had one of the loadcell cables going into the amp pulled out.
measured the impedances between legs of the loadcell bridge for td12,td8. Load cells look ok.
110602
repaired
110606

8
lightning strike.
  • little star had to be repaired
  • dac card, signal conditioning. Had blown chips
  • uio1 had to be replaced
  • encoder changed
  • amplifier replaced.
  • blown chip on signal conditioning --> strike went through amplifier up to the signal conditioning..
  • accidently installed the old software for the roc 424  encoder rather than roq425.  The sequence was:
    • jack and encoder at 11.2994
    • failure (jack remains at 11.2994 since emg stop on)
    • fixed little star. installed wrong encoder software.
    • old encoder reads 12.776 .. assumed old encoder was bad. jack still at 11.2994
    • installed new encoder, moved encoder until it read 11.2994 (without moving jack).
    • moved jack and encoder up, down. At some point realized that wrong software was installed.
    • moved jack and encoder (with wrong software) to 11.2994. Hopefully this is the same jack position as the failure.
    • installed correct encoder software. Encoder now read 12.776 (with jack still at 11.2994)
    • moved encoder (jack stationary) till encoder read 11.2994. Hopefully we now have the encoder and jack both at 11.2994.
      • We are assuming that with the wrong software installed, returning the encoder to the same encoder reading returned the jack to the same jack position.
110425

12
recalibrated td12 load cells.
ldcell1 : was pretty close.. updated balance and gain.
ldcell2: dc offset of -.237 volts. adjusted balance and only got -.118 mv left it this way...
             didn't modify the   gain.
101014

12
td was oscillating. replaced encoder and started to work ok.
100919

12
td 12 position has been jumping to 24 inches a lot. replaced cable bus to encoder board on 17sep10. But problem persisted. Today replaced encoder interface board.
100911

12,8
recalibrated load cells. td8 AtoD board still had R8 resistor in (i cut it).
100823

4
load cells adjusted now ok, cut R8 resistor on A/D and adjusted amps.
20aug10 adusted but a/d board still had resistor R8 in .
Still need to work on Td12, td8.
100813

12,4,8
lightening damages all 3:
td12: dac, encoder board
td4:   little star, dac,adc, 24 volt supply
td8: little star, DI1, and probably more.
tensions are suspect...
100326

4
replaced amplifier on td4. Was giving motor output fault.
100203

12
found broken beam where td 12 backstays connect to platform (left side looking from north). Not sure when it broke.
090919

12
load cell 12-2 failed. replaced with new load cell 06oct09
090204

12
readjusted the aluminum plate that holds the limit switches. It had been mounted about 2.5 inches off when the jack was replaced. We had no lower limits (0 inches) since the jack replacement. limits are now ok.
090109

12
replaced jack td 12
080623

12
load cell 2 replaced with spare that had come back from being calibrated.
080619

4
transmission replaced. loadcell recalibrated.
080616

12
Found that load cell amplifier was causing trouble with 5 volts supply. After disconnecting amplifier the analog inputs read ok and the encoder stopped jumping.
080608

12
encoder card on td 12 blown out by lightning.
080529

4
replaced jack on tiedown 4.
080407

8
td 8 stopped moving at 19.57. commanded velocity remained 0.
080407

12,4,8
the torque readback has occasional level shifts of up to 55 a/d counts.
080312

12,4,8
recalibrated all 3 load cells. td 12 now has hectors amp adjustement board (it was the last to not have it).
080228

12
replaced transmission. Had a loud knocking noise ...
071011

-
after 2nd set of az swings updated the reference position to be:
15.0718,15.3196,15.0926
071010

-
updated the tiedown reference position. After azswing updated the reference position to be
15.1140,15.2028,15.1670
071209

4,8
moved the nuts on td4 down .5 inches, td8 nut up 3.25 inches to equalize the tiedown extensions.
071004

12
motor faults. replaced motor.
070328
4,8 New amplifiers with hector's calibration boards installed on ldcells. Recalibrated ld cells.
070206
4 ldcells recalibrated
061114
12 ldcell 1 replaced. all load cells recalibrated 
061111
12 ldcell 1 has bad resistance on 1 of the legs of the wheatstone bridge.  ldcell2 is driving both inputs now. 
061108
12 recalibrated both load cells. load cell 1 had water in the cover at the load cell.
061101
4,8 recalibrated load cells.
061028
12 replaced ldcell2 amplifier that was blown out on 061014.
061014
8
12
lightning strike blew out td8. Had to change little star, some of the i/o boards, and some of the analog signal conditioning boards.
td12 blew out ldcell2 amplifier. This is one of the new amplifiers.
060908
12 new online ups installed in td 12. 
060831
12 forgot to swap the pots when the amps were swapped yesterday. corrected at 9:00 am
060830
12 load cell cable before the amplifiers back to normal order. ldc1-> output ldc1 now.
060824
12 load cell cables at load cell put back to normal order
load Cell cable before amplifiers switched. ( in the morning).
060823
12 switched the load cell cables as they leave the load cells.
060822
12 cleaned the load cell cables where they connect to the load cells. They were pretty dirty (full of corrosion)
060814
12 swapped td12 ldcells back to normal position. They had been crossed at the analog input since 060719. Soldcell1 is again ldcell 1. Found the jumping problem was the db25 connector at box. Cleaning this fixed it. This td has no filters.
060812
12 td12 encoder reading failed around noon after a power dip. The encoder always read 9 inches and did not report an error. When requested to move to a higher or lower position (than 9 ") the tiedown would move until it hit the limit switch (since it was looking at the encoder to see when it got there). I manually moved td 12 to 12" using the tape measure.  I moved the other two tiedowns to the correct position when td 12 is at 12". This position is close to where the tiedowns were the previous nite.
I checked the voltage on the encoder board and it was ok (5 volts). On 060814 the problem had fixed itself.
060719
12 swapped td load cell inputs to a/d signal conditioning board to see if problem with ldcell 1 jumps to load cell 2.
060628
12 td 12 ldcell 1 started to fail on 28jun06. jumped to 40 kips and then back. Remained bad thru at least 17jul06 (date of this entry).
060612
8
12,4,8
td 8 ldcell #1 amp changed.
load cells recalibrated.
td 12 ld cell 2  amplifier switched to new amplifier
060526
12 analog output values wrong (velocity feedback not eq  to requested). Replaced analog output signal conditioning brd. The same problem occurred for a short thime on 21may06. 
 The zero vel offset for the dacs jumped back up to 2016. There is something funny going on when we replace the analog output board for td12. The offset either sits around 2020 or 2000. Are we mixing the old and new boards when we replace them ??? see /home/phil/vw/datatk/pnt/ttt/tie/tieParms.log (at the bottom) for the history of the offsets.
060426
12 analog output board fails. replaced.
060420
12 digital i/o output was causing the master contactor to not close reliably (more info)
060306
12 encoder replaced. Also noted that the cable that encoder plugged into may have been dirty . cleaned with water displacer.
060303
12 replaced encoder interface board.
060302
12 td12 releases because of bad enc reading.
051027
all installed db25 connector filters for load cell inputs  and pcu .
051026
td12 replaced analog output board and analog output signal condioning board.
051014
td4 replaced amplifier. It went bad about a week ago. Recalibrated amp1 , amp2 of td4
050916
td8 the analog inputs  from td 8 had trouble during the morning of 16sep05. The plots show all of the  analog inputs vs time for the morning of 16sep05 (.ps) (.pdf) (black td12,red td4,green td8). The first 6 (+24volts, -24volts, brake24volts, ldcell1, ldcell2, and encoder 5 volts all jumped together. At 2:50 they jumped to the maxium value for a few minutes. From 3:30 to 4:00 and again at 6:10 they where at the maximum value. The velocity feedback and master amp speed monitor did not have this problem (the short spikes they have is rfi in the cabinet). The master amp current looks like it jumped up at 2:50 but not at the later times. Need to find out what these inputs have in common.
050914
td4 Load cells out of calibration (again). ldcell1 amp had dc offset of 1.5 volts. Both load cells had small oscillations in the output (with no input load). Replaced both load cell amplifiers.
050904
td12 db25 filter for pcu cable shorted out 24 volt supply. removed filter.
050902
td8 both ldcell amplifiers 8-1, 8-2 had their excitation voltage at 0. replaced both load cell amplifiers then recalibrated. This had happened a while back (probably on 25aug05.
removed the db25 filter for loadcell 8-1. it was bad (need to send away for more filters).
050902
td4 both load cell amplifiers have excitation voltage of 0. need to be fixed..
050825
td8 lightning.. replaced little star
050825
td12 lightning .. replaced encoder interface board, signal conditioning for dac, analog input fuse for input 3?
050728
td12 trouble with td 12. replaced little star. releasing.
050625
td12 lightning trouble. See writeup. Ended up replacing little start, amp.
050317
td12 replaced jack td 12. looked at prelim encoder position. moved by about .2 inches.
050317
td12 replaced encoder board on td12. The encoder value had jumped from 13 to 24 inches at around 9am.
050225
td4 td 4 jack breaks. looks like bad ball bearing. replaced jack 28feb05. reset limit switches, encoder.
050131
td12 changed filter ldcell1,2. then the values were fine. Must have been water in the filter.
050126
td12 td12 ldcell 2 value value jumped to a lower reading. this means that both load cells are reading low. 
050124
ts12 td12 ldcel 1 dropped about 10kips in 45 minutes. 
050117
td4 replaced filter on ldc1 cable. it was bad..
recalibrated all ldcells for the 3 tiedowns
041116
td4 still trouble with td 4. occasionally when command from stop, td 4 would 
1.release the brakes 
2. wait a few seconds 
3. go max speed.. this caused and overspeed fault . 
replaced amplifier. Still had trouble. 
replaced a number of little stars, had trouble with parameters. Loaded ok but when we 
cycled power on LS, they went back to the original values. 
CB2,3 tripped again. Had to disable the fan on the box. Then it worked ok. 
Tried another little star and this time it worked. Left the box off till we could 
get the box fan working.
041112
td4 jack released. kolmorgan power supply failed...replaced 15nov04. It had a short. 
also changed the dac board and signal conditioning.
040510
td4 loud noise from the jack. replaced the jack. This required dismantling the limit switches, encoder and then replaced after new jack installed. reference may have moved some. (our tape measurement were not longer valid since new jack).
040317
td8 safety relay problem was water in liftoff switches. Only had 15 volts rather then 24. Cleaned the switch and it started working.
040317
td8 db25 feed thru filter for pc. Replaced it and the pcu started working.
040315
pcu pcu was not working . brought it up to the lab to check it out. (problem was feed thru filter td8)
040315
8 safety relay was not closing with result pulse. after cycling it manually a few times it started to work. Switched pcu relay with safety relay. problem was feed water in switch. ( see 040317)
040315
4 encoder jumping to 24". started 4th march, 12 march shut off. 15mar replaced encoder intf board.
030930
mod4 changed little star td4 (battery was low). updated parameters in little star.
030806
mod 
8
recalibrated loadcells on td 8
030806
mod 
4
rotated mounting bracket for limit switches to face away from the box. Old tape  measurements no longer valid for td4.
030730
mod 
8
final  lower limit td 8 reset to be at encoder position 1.19". Ran the jack all the way to 0" encoder value and still did not hit end of jack (although it is probably pretty close).
030728
Note The limit switches were changed to a newer type that did not leak (i'm not sure what date). It looks like they were mounted in the same position as the old switches even though the mechanical layout is different. The jack positions for where the limits triggered are now different. For td8, the jack runs all the way to the end of the jack and the limit switch has still not tripped. This is what caused the limits to not trip on 030728. All of the limit switches need to be repositioned so they are at the correct jack positions.
030728
Fail lightning afternoon.
 
8 blew out dac. Td went all the way up. Final limit switch did not trip.
Replaced dac. 
Final limit switch has moved by more than 1 inch. Probably not the encoder since we actually ran into the end of the jack without hitting the limit.
 
4 Encoder board blown. Replaced board. Encoder before problem was 13.301. After it was 12.6?? Someone must have moved the td while board was bad. Since abs encoder was ok, position is probably ok
031111
Fail 
4
td4 drive disabled itself even though program thought it was still running. 
Luis found that the voltage regulator for the dac was shorted. Replaced dac.
021003
mod 
all
5 volt supply for dacs is now coming from a separate power supply. It used to come from the dc to dc converter on the encoder board. Change made to t12. Need to also update the others. Need to move the 5 volt encoder voltage monitoring line from the dac to the a/d.
021002
Fail lightning
 
12 blew out encoder interface. replaced. The kips were reading 81,82 until the encoder board was replaced. The lcd display on 12 was also acting up.
 
8 blew out encoder interface board. replaced. 
Recalibrated td8 load cells.
020923 

 


fail 
8
Axis fault always on. tiedown was oscillating. This can be from encoder init error or master contactor did not close after 20 cycles. This encoder interface board had been installed 10sep02. Replacing the encoder interface board fixed the problem.
Updated the zero velocity offsets for td12 and td8.
020910
fail lightning struck platform.
 
12 signal conditioning bad. Dac may also have been blown. Swapping encoder interface board screwed up dac or signal conditioning again.. This was an old encoder interface board. roc 424.
 
8 encoder bad. Dac or signal conditioning also bad.
020815
fail power dip or lightning??
 
12 dac output reading -2048 counts. cycling power fixed it.
 
4 dac signal conditioning board blew. Outputting -24 volts.
 
8 encoder reading 24 inches. encoder interface replaced with new encoder board/pal (425q pals.
020424 
 
 

 


fail 

 

 

power dip 
td8 not responding. +24 pwr supply has 3.5 volts. Removing little star fixed it so little start was shorting out pwr supply. Put in new little star, didn't work. Put in 2nd little start. Worked but had to update parameters. After updating parameters, would only move up , not down. Maxpos was 0 so i set it to 24000. Next day rebooted and it worked ok. Looks like for some parameters you need a reboot to take affect.
011203
fail8 lightning . td 8 dead. Short in +24 volt power supply.
011101
fail8 little start digital input lspin09 bad. opto-isolator on signal conditioning board shot. This caused the master contactor to always read closed, even when it was open.
011101
all Tiedowns back on the air. They'd been off since 23sep01
011101 
 
 

 


fail 

 

 

amplifier for td8 amp had missing connections on a capacitor, resistor, and a blown capacitor. Replaced amp with repaired amp from kolmorgan. This amp had speed offset for the azimuth (it was an az motor). On td this caused and over speed. The dac output was 8 volts and the speed output (of amp) was over 11 volts. Tried adjusting a pot to bring the speed voltage to 0 but it looks like it screwed up the gain. Installed another amp and it worked ok. 
011031 

 


fail8 

 

load cell td8 amp1 jumped 1.2 volts in dc offset. (ldcell:#90407). Switched cables at output of load cells and still had the problem. Removed db25 connector at case and it worked ok. Replaced connector (these connectors have built in filters). Measured the resistance for the wheatstone bridge in td8,4 load cells. records in manual so you can easily check if the db25 connector is blown (it changes the resistance).
011030
mod Td 8,td4 now have roq424/425 encoders. They work fine. Td 12 has old roq424 encoder. it jumps in the 3rd hex digit from the right.
Td8 A/D card jumper was not cut. did it today.
011029
mod Replaced ldcell amp td8 #1. Recalibrated (but the a/d resistor had not been cut.
010923
fail lightning. tiedowns remain down until 01nov01
 
12 
 
 
 

 

encoder moving in wrong direction. Changed encoder, started to move and then little start died. changed little start. Luis had measured the position at 14.7 inches. So moved td to luis's tape measurement and then set the encoder to 14.7 inches. Then went to the limits and got: 
expected     measured
1.43              1.8473        .4 inches off 
22.29          22.28             .01 inches off
 
4 encoder reads 24 inches. resetting td 4 seemed to fix it , but moving 4 then made a terrible noise. Dac is shot, red light is off. Replaced dac and switched to single ended mode on signal conditioning board (no jumper and 10k resistor).
 
8 encoder reads 24 inches. cycling power it then reads 0 inches. 
Lifted bus cable to dac,di1,di2 and a/d started working. looks like dac is screwing up bus. Replaced dac, signal conditioning. still troubles. looks like encoder board is messing up the bus (need to build some more encoder boards for the new encoders).

Major structural changes
date
td
description
170901
8
tiedown jack replaced
170827
12
tiedown jack replaced
170825
4
tiedown jack replaced.
1601xx
8
load cell 2 replaced. damaged during hf run back in nov15
120423
12
tiedown jack replaced
120516
8
jack replaced? more info
090109
12
tiedown jack replaced
080619
4
transmission replaced
080529
4
tiedown jack replaced
080228
12
transmission replaced. had a loud knocking sound
050605
8
according to felipe soberal
050317
12
tiedown jack replaced
050225
4
tiedown jack replaced.

<- page up
home_~phil